Bailey, Inc., is considering buying a new gang punch that would allow them to produce circuit boards more efficiently. The punch has a first cost of $160,000 and a useful life of 15 years. At the end of its useful life, the punch has no salvage value. Labor costs would increase $5,500 per year using the gang punch, but raw material costs would decrease $16,500 per year. MARR is 5%/year. 


What is the internal rate of return of this investment? 


What is the decision rule for judging the attractiveness of investments based on internal rate of return? 


Should Bailey buy the gang punch? 
Answer #1

Net annual savings = decrease in raw material cost-icrease in labor cost for using punch =16500-5500= \$ 11000

Present value of an annuity is given as

\begin{aligned} & P\left[\frac{1-(1+r)^{-n}}{r}\right] \\ P=& \text { Periodic Payment } \\ r=& \text { rate per period } \end{aligned}

n= number of periods

n=15

Let IRR =r

For IRR, NPV=0

i.e. initial cost =\mathrm{PV} of annuity

160000=11000^{*}\left(1-(1+r)^{\wedge}-15\right) / r

\left(1-(1+r)^{\wedge}-15\right) / r=160 / 11

By hit and trial ,method, we get r=0.4 \%

Decision rule for project approval is MARR > IRR

No, Bailey should not buy this punch as IRR

answered by: Coatorly